A board 10ft long is cut into two pieces. Thgree times the length of the shorter pice is twice the length of the longer piece. Find the length of each piece.
Answer by rcmcc(152)

we can solve this question by making a two linear equations to solve for the shortest piece.
lets make the shortest piece x
the middle piece y
since the longest board is 3 times larger than the smallest board we can express it as 3x
3x=2y
we can make the standard equation by adding all the x's and y to equal 10
3x+x+y=10
by using the substitution method you can solve for x by isolating y in the first equation and substituting it in for x
(3/2)x=y
4x+(3/2)x=10
5.5X=10
divide both sided by 5.5 to isolate x
x=(20/11)
So the largest piece is (60/11)ths and the middle piece is (30/11)ths.
so your answers are
smallest peice=(20/11)ft.
middle piece=(30/11)ft.
largest piece=(60/11)ft.